# Atomic sheaves on hyper-Kähler manifolds via Bridgeland moduli spaces

### v2

@@ 7.2.1. LLSvS eightfolds
Let $X$ be a cubic fourfold not containing a plane and $j\colon Y\hookrightarrow X$ be a smooth hyperplane section. Let $M_{3}(X)$ be the Hilbert scheme of generalized twisted cubics on $X$, which is a smooth projective tenfold by [48, Theorem 4.7].
@@
On the other hand, we have the moduli interpretation $\mathsf{Z}_{X}\cong M^{X}_{\sigma_{X}}(2,1)$ (see [50]) and $\mathsf{Z}_{Y}\cong M^{Y}_{\sigma_{Y}}(2,1)$ (see [3] or [25, Appendix B.2]). In particular, if we denote by $i\colon\mathsf{Z}_{Y}\hookrightarrow\mathsf{Z}_{X}$ the Lagrangian embedding constructed in [70], the moduli construction is shown to be compatible with $i$ by [25, Corollary B.7], as the following commutative diagram
@@
is of rank one when $X$ is non-Hodge-special. Furthermore, a deformation argument using Theorem 3.2 and 3.1 shows that this holds for a general $X$ as well. Subsequently, based on Theorem 4.10 and Theorem 6.7, one might speculate that $\mathsf{Z}_{Y}\subset\mathsf{Z}_{X}$ is an atomic Lagrangian submanifold. However, the following result demonstrates that this is not the case.
@@
    $i_{*}\mathcal{L}\in\operatorname{\mathrm{Coh}}(F(X))$ is a stable atomic sheaf for any line bundle $\mathcal{L}$ on $\mathsf{Z}_{Y}$ satisfying
@@
induced by the Stein factorization of the natural morphism $M_{3}(X)\rightarrow\operatorname{Gr}(4,V_{6})$, which maps a twisted cubic to its generating $\mathbb{P}^{3}$. Moreover, $\mathsf{Z}^{\prime}_{X}$ is smooth by [48, Theorem 4.7]. Similarly, there exists a smooth subvariety $\mathsf{Z}^{\prime}_{Y}\subset\mathsf{Z}^{\prime}_{X}$ induced by $M_{3}(Y)\subset M_{3}(X)$, and $f_{X}$ induces a morphism $f_{Y}\colon\mathsf{Z}^{\prime}_{Y}\to\operatorname{Gr}(4,V_{5})$, which maps a twisted cubic on $Y$ to its generating $\mathbb{P}^{3}$ as well. Therefore, it is straightforward to check that there is a Cartesian diagram
@@
According to [48, Theorem 4.11], there is an embedding $X\hookrightarrow\mathsf{Z}_{X}$ such that a point $[C]\in M_{3}(X)$ maps to $p\in X\subset\mathsf{Z}_{X}$ if and only if $C$ has $p$ as an embedded point. Moreover, the blow-up of $\mathsf{Z}_{X}$ along $X$ is $\mathsf{Z}_{X}^{\prime}$. We denote by $D\subset\mathsf{Z}_{X}^{\prime}$ the exceptional divisor, and let $\phi\colon\mathsf{Z}_{X}^{\prime}\to\mathsf{Z}_{X}$ be the blow-up morphism. As any two twisted cubics with embedded points in $Y$ have the same embedded point if and only if they span the same $\mathbb{P}^{3}$, we have an identification $\mathsf{Z}_{Y}^{\prime}=\mathsf{Z}_{Y}$. Then we obtain a commutative diagram
@@
By [48, pp.38], there exists an ample line bundle $L$ on $\mathsf{Z}_{X}$ such that we have an isomorphism $\operatorname{\mathcal{O}}_{\mathsf{Z}_{X}^{\prime}}(D)\otimes f^{*}_{X}\operatorname{\mathcal{O}}_{\operatorname{Gr}(4,V_{6})}(1)\cong\phi^{*}L$. Moreover, according to [48, pp.37], we have $\omega_{\mathsf{Z}^{\prime}_{X}}\cong\operatorname{\mathcal{O}}_{\mathsf{Z}_{X}^{\prime}}(3D)$, which implies
@@
then $D|_{\mathsf{Z}_{Y}}=0\in\mathrm{H}^{2}(\mathsf{Z}_{Y},\mathbb{Q})$. However, this is impossible since $\mathcal{O}_{Y}(D|_{Y})\cong\operatorname{\mathcal{O}}_{Y}(-1)$ by [48, Proposition 4.5]. Therefore, we see $c_{1}(\mathsf{Z}_{Y})\notin\operatorname{im}(i^{*})$ and $\mathsf{Z}_{Y}\subset\mathsf{Z}_{X}$ is not atomic by Theorem 4.10. This finishes the proof of (1).
@@
Applying the argument in Theorem 1.2 to the atomic sheaf in Proposition 7.3(2), we have the following result.
@@ 7.2.2. Fano variety of lines
Besides this construction, the authors of [35] consider another Lagrangian surface of $F(X)$ arising from a cubic fivefold. Let $Z$ be a cubic fivefold, where $X$ is a hyperplane section of $Z$. We consider the Fano variety of planes of $Z$, denoted by $F_{2}(Z)$. When $X$ and $Z$ are general, $F_{2}(Z)$ is a smooth surface and there is a morphism $i\colon F_{2}(Z)\to F(X)$ such that $F_{2}(Z)$ is a connected immersed Lagrangian submanifold of $F(X)$ (cf. [35, Proposition 4]). Using a standard deformation argument of relative Hilbert schemes, one can show that $i^{*}\colon\mathrm{H}^{2}(F(X),\mathbb{Q})\to\mathrm{H}^{2}(F_{2}(Z),\mathbb{Q})$ is of rank one, i.e. these Lagrangians deform in codimension one. However, we have the following result.
@@
    $i_{*}\mathcal{L}\in\operatorname{\mathrm{Coh}}(F(X))$ is a stable atomic sheaf for any line bundle $\mathcal{L}$ on $F_{2}(Z)$ satisfying $c_{1}(\mathcal{L})=7(h_{2}-h_{1})\in\mathrm{H}^{2}(F_{2}(Z),\mathbb{Q})$, where $h_{1}$ is the pull-back of the Plücker class via the natural morphism $F_{2}(Z)\to\operatorname{Gr}(3,7)$ and $h_{2}$ is the pull-back of the Plücker class on $F(X)$ via $i$.
@@
We assume that $X=Z\cap\mathbb{P}(V_{6})\subset\mathbb{P}(V_{7})$. Let $\mathcal{T}^{\prime}$ be the pull-back of the rank two tautological subbundle on $\operatorname{Gr}(2,V_{6})$ via the natural embedding $F(X)\hookrightarrow\operatorname{Gr}(2,V_{6})$ and $\mathcal{W}:=V_{6}\otimes\operatorname{\mathcal{O}}_{F(X)}$. In [35, Lemma 6], it is shown that $i$ can be decomposed as
@@
If we denote by $\alpha:=c_{1}(\operatorname{\mathcal{O}}_{\mathbb{P}}(1))$ and $\beta:=c_{1}(\mathcal{T}^{\vee})$, then we get $c_{1}(\mathcal{K})=10\alpha+3\beta$ and $c_{1}(\mathbb{P})=4\alpha+\beta$. Hence, $c_{1}(F_{2}(Z))=14t^{*}\alpha+4t^{*}\beta$. However, as $h_{2}=t^{*}\beta$ is the pull-back of the class of the Plücker line bundle on $F(X)$, we have $0\neq t^{*}\beta\in\operatorname{im}(i^{*})$ and $\operatorname{im}(i^{*})$ is spanned by $t^{*}\beta$. Thus, as $c_{1}(F_{2}(Z))\in\operatorname{im}(i^{*})$, we have $t^{*}\alpha=0\in\mathrm{H}^{2}(F_{2}(Z),\mathbb{Q})$. Since the fundamental class of $F_{2}(Z)$ in $\mathbb{P}$ is $63\alpha^{4}c_{2}(\mathcal{T}^{\vee})$, we obtain $063\alpha^{5}c_{2}(\mathcal{T}^{\vee})=\alpha.t_{*}[F_{2}(Z)]=t_{*}t^{*}\alpha=$ 0 . However, as computed in [35, Lemma 6], we have $\alpha^{5}=-\beta.\alpha^{4}$. This implies that
@@
Applying the argument in Theorem 1.2 to the atomic sheaf in Proposition 7.5(2), we have the following result.

### v3

@@ 7.2.1. LLSvS eightfolds
Let $X$ be a cubic fourfold not containing a plane and $j\colon Y\hookrightarrow X$ be a smooth hyperplane section. Let $M_{3}(X)$ be the Hilbert scheme of generalized twisted cubics on $X$, which is a smooth projective tenfold by [52, Theorem 4.7].
@@
On the other hand, we have the moduli interpretation $\mathsf{Z}_{X}\cong M^{X}_{\sigma_{X}}(2,1)$ (see [55]) and $\mathsf{Z}_{Y}\cong M^{Y}_{\sigma_{Y}}(2,1)$ (see [3] or [27, Appendix B.2]). In particular, if we denote by $i\colon\mathsf{Z}_{Y}\hookrightarrow\mathsf{Z}_{X}$ the Lagrangian embedding constructed in [75], the moduli construction is shown to be compatible with $i$ by [27, Corollary B.7], as the following commutative diagram
@@
is of rank one when $X$ is non-Hodge-special and $\ker(i^{*})=h^{\perp}$, where $h$ is the natural polarization on $\mathsf{Z}_{X}\cong M^{X}_{\sigma_{X}}(2,1)$. Furthermore, a deformation argument using Theorem 3.2 and 3.1 shows that this holds for a general $X$ as well. Subsequently, based on Theorem 4.10 and Theorem 6.7, one might speculate that $\mathsf{Z}_{Y}\subset\mathsf{Z}_{X}$ is an atomic Lagrangian submanifold. However, the following result demonstrates that this is not the case.
@@
    $i_{*}\mathcal{L}\in\operatorname{\mathrm{Coh}}(\mathsf{Z}_{X})$ is a stable atomic sheaf for any line bundle $\mathcal{L}$ on $\mathsf{Z}_{Y}$ satisfying
@@
induced by the Stein factorization of the natural morphism $M_{3}(X)\rightarrow\operatorname{Gr}(4,V_{6})$, which maps a twisted cubic to its generating $\mathbb{P}^{3}$. Moreover, $\mathsf{Z}^{\prime}_{X}$ is smooth by [52, Theorem 4.7]. Similarly, there exists a smooth subvariety $\mathsf{Z}^{\prime}_{Y}\subset\mathsf{Z}^{\prime}_{X}$ induced by $M_{3}(Y)\subset M_{3}(X)$, and $f_{X}$ induces a morphism $f_{Y}\colon\mathsf{Z}^{\prime}_{Y}\to\operatorname{Gr}(4,V_{5})$, which maps a twisted cubic on $Y$ to its generating $\mathbb{P}^{3}$ as well. Therefore, it is straightforward to check that there is a Cartesian diagram
@@
According to [52, Theorem 4.11], there is an embedding $X\hookrightarrow\mathsf{Z}_{X}$ such that a point $[C]\in M_{3}(X)$ maps to $p\in X\subset\mathsf{Z}_{X}$ if and only if $C$ has $p$ as an embedded point. Moreover, the blow-up of $\mathsf{Z}_{X}$ along $X$ is $\mathsf{Z}_{X}^{\prime}$. We denote by $D\subset\mathsf{Z}_{X}^{\prime}$ the exceptional divisor, and let $\phi\colon\mathsf{Z}_{X}^{\prime}\to\mathsf{Z}_{X}$ be the blow-up morphism. As any two twisted cubics with embedded points in $Y$ have the same embedded point if and only if they span the same $\mathbb{P}^{3}$, we have an identification $\mathsf{Z}_{Y}^{\prime}=\mathsf{Z}_{Y}$. Then we obtain a commutative diagram
@@
By [52, pp.38], there exists an ample line bundle $L$ on $\mathsf{Z}_{X}$ such that we have an isomorphism $\operatorname{\mathcal{O}}_{\mathsf{Z}_{X}^{\prime}}(D)\otimes f^{*}_{X}\operatorname{\mathcal{O}}_{\operatorname{Gr}(4,V_{6})}(1)\cong\phi^{*}L$. Moreover, according to [52, pp.37], we have $\omega_{\mathsf{Z}^{\prime}_{X}}\cong\operatorname{\mathcal{O}}_{\mathsf{Z}_{X}^{\prime}}(3D)$, which implies
@@
then $D|_{\mathsf{Z}_{Y}}\in\operatorname{im}(i^{*})$ is either ample, zero, or anti-ample. As $\mathcal{O}_{Y}(D|_{Y})\cong\operatorname{\mathcal{O}}_{Y}(-1)$ by [52, Proposition 4.5], we know that $-D|_{\mathsf{Z}_{Y}}$ is ample. In particular, we obtain
$\mathrm{H}^{0}(\mathsf{Z}_{Y},\operatorname{\mathcal{O}}_{\mathsf{Z}_{Y}}(D|_{\mathsf{Z}_{Y}}))=\mathrm{H}^{4}(\mathsf{Z}_{Y},\omega_{\mathsf{Z}_{Y}}(-D|_{\mathsf{Z}_{Y}}))=0$
by Kodaira vanishing theorem, which leads to a contradiction since $Y=D\cap\mathsf{Z}_{Y}\in|\operatorname{\mathcal{O}}_{\mathsf{Z}_{Y}}(D|_{\mathsf{Z}_{Y}})|$. Therefore, we conclude that $c_{1}(\mathsf{Z}_{Y})\notin\operatorname{im}(i^{*})$ and $\mathsf{Z}_{Y}\subset\mathsf{Z}_{X}$ is not atomic by Theorem 4.10. This completes the proof of (1).
@@
Applying the argument in Theorem 1.2 to the atomic sheaf in Proposition 7.3, we have the following result.
@@
Proof.
Since $i\colon\mathsf{Z}_{Y}\hookrightarrow\mathsf{Z}_{X}$ is a closed embedding that realizes $\mathsf{Z}_{Y}$ as the zero locus of a regular section of a rank $4$ bundle on $\mathsf{Z}_{X}$ as in the proof of Proposition 7.3, we have
$\operatorname{Ext}^{1}_{\mathsf{Z}_{X}}(i_{*}\mathcal{L},i_{*}\mathcal{L})\cong\mathrm{H}^{1}(\mathsf{Z}_{Y},\mathbb{C})\cong\mathbb{C}^{10},$
where $\mathcal{L}$ is defined in Proposition 7.3. Alternatively, we can also use Lemma 5.12, where the family $L_{V}\to V$ is constructed from the relative Bridgeland moduli space of the smooth universal hyperplane in $X$.
We know that the locus of $\mathsf{Z}_{X}$ form an open dense subset of the moduli space of projective hyper-Kähler manifold of degree $3$ and divisibility $2$ (cf. [52]). Then by the Torelli Theorem [35, Theorem 8.1], we can take $X$ such that $\mathrm{H}^{1,1}_{\mathbb{Z}}(\mathsf{Z}_{X})=\langle h,f\rangle$, where $h$ is the natural polarization and $f$ induces a Lagrangian fibration $\mathsf{Z}_{X}\to\mathbb{P}^{4}$ such that $q(h,f)=d$ for a suitable positive integer $d$. Moreover, we have $f^{4}.i_{*}[\mathsf{Z}_{Y}]\neq 0$ by $\ker(i^{*})=h^{\perp}$. Hence, $\mathsf{Z}_{Y}$ is finite over $\mathbb{P}^{4}$ as in Theorem 1.2, since $\operatorname{im}(i^{*})$ is spanned by an ample class. Now, the rest of the argument in the proof of Theorem 1.2 works without any change. ∎
@@ 7.2.2. Fano variety of lines
Besides this construction, the authors of [39] consider another Lagrangian surface of $F(X)$ arising from a cubic fivefold. Let $Z$ be a cubic fivefold, where $X$ is a hyperplane section of $Z$. We consider the Fano variety of planes of $Z$, denoted by $F_{2}(Z)$. When $X$ and $Z$ are general, $F_{2}(Z)$ is a smooth surface and there is a morphism $i\colon F_{2}(Z)\to F(X)$ such that $F_{2}(Z)$ is a connected immersed Lagrangian submanifold of $F(X)$ (cf. [39, Proposition 4]). Using a standard deformation argument of relative Hilbert schemes, one can show that $i^{*}\colon\mathrm{H}^{2}(F(X),\mathbb{Q})\to\mathrm{H}^{2}(F_{2}(Z),\mathbb{Q})$ is of rank one and $\ker(i^{*})=h^{\perp}$, where $h$ is the natural Plücker polarization on $F(X)$. However, we have the following result.
@@
    $i_{*}\mathcal{L}\in\operatorname{\mathrm{Coh}}(F(X))$ is a stable atomic sheaf for any line bundle $\mathcal{L}$ on $F_{2}(Z)$ satisfying $c_{1}(\mathcal{L})=7(h_{2}-h_{1})\in\mathrm{H}^{2}(F_{2}(Z),\mathbb{Q})$, where $h_{1}$ is the pull-back of the Plücker class via the natural morphism $F_{2}(Z)\to\operatorname{Gr}(3,7)$ and $h_{2}$ is the pull-back of the Plücker class $h$ on $F(X)$ via $i$.
@@
We assume that $X=Z\cap\mathbb{P}(V_{6})\subset\mathbb{P}(V_{7})$. Let $\mathcal{T}^{\prime}$ be the pull-back of the rank two tautological subbundle on $\operatorname{Gr}(2,V_{6})$ via the natural embedding $F(X)\hookrightarrow\operatorname{Gr}(2,V_{6})$ and $\mathcal{W}:=V_{6}\otimes\operatorname{\mathcal{O}}_{F(X)}$. In [39, Lemma 6], it is shown that $i$ can be decomposed as
@@
If we denote by $\alpha:=c_{1}(\operatorname{\mathcal{O}}_{\mathbb{P}}(1))$ and $\beta:=c_{1}(\mathcal{T}^{\vee})$, then we get $c_{1}(\mathcal{K})=10\alpha+3\beta$ and $c_{1}(\mathbb{P})=4\alpha+\beta$. Hence, $c_{1}(F_{2}(Z))=14t^{*}\alpha+4t^{*}\beta$. However, as $h_{2}=t^{*}\beta$ is the pull-back of the class of the Plücker line bundle on $F(X)$, we have $0\neq t^{*}\beta\in\operatorname{im}(i^{*})$ and $\operatorname{im}(i^{*})$ is spanned by $t^{*}\beta$. Thus, as $c_{1}(F_{2}(Z))\in\operatorname{im}(i^{*})$, we have $t^{*}\alpha=0\in\mathrm{H}^{2}(F_{2}(Z),\mathbb{Q})$. Since the fundamental class of $F_{2}(Z)$ in $\mathbb{P}$ is $63\alpha^{4}c_{2}(\mathcal{T}^{\vee})$, we obtain $063\alpha^{5}c_{2}(\mathcal{T}^{\vee})=\alpha.t_{*}[F_{2}(Z)]=t_{*}t^{*}\alpha=$ 0 . However, as computed in [39, Lemma 6], we have $\alpha^{5}=-\beta.\alpha^{4}$. This implies that
@@
Applying the argument in Theorem 1.2 to the atomic sheaf in Proposition 7.5, we have the following result.
@@
Proof.
Since $i\colon F_{2}(Z)\to F(X)$ is a closed embedding away from finitely many points and there exists a nice family of immersed Lagrangians of $F(X)$ as required in Lemma 5.12 (cf. [39]), we have
$\operatorname{Ext}^{1}_{F(X)}(i_{*}\mathcal{L},i_{*}\mathcal{L})\cong\mathrm{H}^{1}(F_{2}(Z),\mathbb{C})\cong\mathbb{C}^{42}$
by applying Lemma 5.12, where $\mathcal{L}$ is defined in Proposition 7.5.
We know that the locus of $F(X)$ form an open dense subset of the moduli space of projective hyper-Kähler manifold of degree $6$ and divisibility $2$ (cf. [4]). Then by the Torelli Theorem [35, Theorem 8.1], we can take $X$ such that $\mathrm{H}^{1,1}_{\mathbb{Z}}(F(X))=\langle h,f\rangle$, where $h$ is the natural Plücker polarization and $f$ induces a Lagrangian fibration $F(X)\to\mathbb{P}^{2}$ such that $q(h,f)=d$ for a suitable positive integer $d$. Moreover, we have $f^{2}.i_{*}[F_{2}(Z)]\neq 0$ by $\ker(i^{*})=h^{\perp}$. Hence, $i(F_{2}(Z))$ is finite over $\mathbb{P}^{2}$ as in Theorem 1.2, since $\operatorname{im}(i^{*})$ is spanned by an ample class. Now, the rest of the argument in the proof of Theorem 1.2 works without any change. ∎